from point A which is 30m from the base of a building B, the angle of elevation to the top of the building C is 56 degrees, and to the top of the flag pole CD is 60 degrees. find the length of the flag pole
Answer by stanbon(75887)

from point A which is 30m from the base of a building B, the angle of elevation to the top of the building C is 56 degrees, and to the top of the flag pole CD is 60 degrees. find the length of the flag pole
------------
Draw the picture:
height of building = 30*tan56
---
height of building+pole = 30*tan(60)
---
height of pole) = 30(tan60-tan56) = 7.4847 meters
=================
Cheers,
Stan H.